# RouteKG: A knowledge graph-based framework for route prediction on road networks

### v3

@@ V-C2 Main Results Analysis
Table III shows a comparison of the accuracy of the different methods in predicting future routes on two real-world datasets under three scenarios. Overall, our proposed RouteKG model achieves optimal accuracy on all metrics. It is observed that for all models, the route-level prediction accuracy is lower than the link-level prediction accuracy. This underscores the importance of modeling the consistency between different road links. Comparing the different models, we can find that Deep Learning-based methods achieve higher accuracy in general. This aligns with empirical findings and is further enhanced with the integration of additional information. For instance, models like NetTraj and RouteKG, which incorporate spatial data, outperform simpler models like RNN and its variants. Remarkably, RouteKG outperforms the NetTraj model and other baselines, even without extra information, which highlights the effectiveness of our approach to integrate KG for future route prediction. Comparing under different experimental settings, intuitively, introducing more Goal information progressively improves overall accuracy. In particular, RouteKG’s prediction accuracy is greatly improved after the gradual incorporation of Goal’s information, and at the same time, it also has an accuracy improvement of about 5.41% in comparison with the optimal baseline under the NoGoal condition, which proves the effectiveness of RouteKG in utilizing Goal information, and its applicability under various conditions. Lastly, comparing different datasets, the prediction accuracy of the Chengdu dataset is better than that of the Shanghai dataset, likely because of the larger data size of the former.
@@ V-D Ablation Analysis
Figure 8 compares the performance of RouteKG with its two ablation variants. Notably, RouteKG w/o rerank removes the Rank Refinement Module. Experimental results show that removing this module significantly reduces prediction performance. This suggests the interconnected nature of link choices, emphasizing the need for a module to model route consistency and choice correlation. This highlights the module’s indispensability. Remarkably, even without reranking, RouteKG still outperforms most benchmark methods, particularly in the top 5 and top 10 predictions. This demonstrates RouteKG’s efficacy in identifying potential future routes, reinforcing the importance of integrating the ranking refinement module for enhancing top-1 predictions.
@@ V-E Sensitivity Analysis
We perform a sensitivity analysis to assess the robustness and reliability of our model under various parameter settings.
@@
We investigate the model’s performance under varying lengths of future routes to be predicted, denoted as $\Gamma^{\prime}$. As depicted in Figure 9, by altering $\Gamma^{\prime}$ from 2 to 8, there is a noticeable trend of declining performance with increasing $\Gamma^{\prime}$ values. This indicates that predicting longer routes becomes progressively challenging due to an expanded candidate space and uncertainty, particularly when lacking Goal information.
@@ V-F Efficiency Analysis
Efficient route prediction in transportation systems is paramount, necessitating prompt response for system operators and road users. To assess the model efficiency, we analyze the inference time of various models in two datasets, to examine model efficiencies. This allows us to identify the most efficient and responsive models, which is crucial for ensuring better user experiences and traffic management.
Figure 10 delineates the inference times across various models. Note that the results from Dijkstra and RCM-BC are omitted due to their overly high inference times. All baselines utilize the Spanning Route algorithm to generate future routes except the Markov model. The Markov model leverages pre-computed transition probabilities to sample and generate the top-$k$ predictions through $k$ iterations.
Figure 10: Inference time per 10K requests for different methods in Chengdu and Shanghai (mean ± std).
@@
TABLE IV: Traffic flow estimation results on Chengdu and Shanghai datasets. (mean ± std)
RouteKG demonstrates remarkable efficiency, achieving average inference times of 598.01ms and 244.47ms for every 10k requests on the Chengdu and Shanghai datasets, respectively, with standard deviations of 1.21ms and 19.35ms. In contrast, the Dijkstra model, based on dynamic programming, takes over 38 seconds, and the RCM-BC model exceeds 1000 seconds, rendering them impractical for real-time systems. Models utilizing the Spanning Route algorithm (e.g., RNN, GRU, LSTM, NetTraj, RouteKG) show superior inference times, with less than 400 milliseconds in Chengdu and 250 milliseconds in Shanghai per 10k requests. RouteKG exhibits a marginally higher inference time, likely attributable to the reranking process. Nonetheless, these results highlight RouteKG’s suitability for real-time traffic applications, where rapid processing is essential.
@@ V-G Case Study: Traffic Flow Estimation
In this section, we conduct a case study on traffic flow estimation to demonstrate the practical use cases of RouteKG, utilizing its potential to generate future routes for accurate traffic flow estimations. This can offer key insights into traffic pattern dynamics and enhance the reliability of traffic flow predictions. Specifically, we adopt a sampling-based method for generating future routes to maximize the utility of the top-$k$ future route predictions. Initially, the top-$k$ predictions are converted to a probability distribution using temperature scaling [78]. Subsequently, we sample from the predicted top-$k$ future routes for each observed trajectory based on their probability distribution. The estimated link-level traffic flows are then obtained by aggregating the number of predicted future routes at the link level. To counter the effects of uncertainty, we iterate the experiments ten times and reported traffic flow estimation results in a mean±std format, focusing solely on the top-10 predictions for simplicity.
The effectiveness of traffic flow estimation with RouteKG is demonstrated using three standard regression metrics: Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and the coefficient of determination ($\mathrm{R}^{2}$), as detailed in Table IV. RouteKG consistently outperforms in all metrics for both datasets, aligning with our main experiment results in Section V-C. Notably, incorporating more Goal information leads to improved accuracy in traffic flow predictions, highlighting the strength of our approach.
In particular, RouteKG’s performance in the NoGoal scenario significantly surpasses the baseline for both datasets, suggesting that our method of estimating moving directions and leveraging KGC is more effective than current state-of-the-art (SOTA) modeling methods. Quantitatively, it reduces MAE, RMSE, and $\mathrm{R}^{2}$ by 34.7%, 43.5%, and 2.6% respectively, compared to the best baseline. Under the GoalD scenario, performance increases notably, indicating potential for future refinement in modeling future directions. Importantly, RouteKG’s enhancements in traffic flow estimation, especially when including the actual future direction, are more significant than those of the baselines. This reaffirms RouteKG’s advanced integration of direction information in the KGC problem. With actual goal information incorporated, RouteKG achieves an MAE of 1, RMSE of 3, and 99.7% in $\mathrm{R}^{2}$, underlining its efficacy and promise for practical applications.
@@ VI Conclusion
RouteKG is evaluated using taxi trajectory data from Chengdu and Shanghai. The evaluation considers three practical scenarios with different levels of goal information availability: NoGoal, GoalD, and Goal. The experiment results show that the proposed RouteKG consistently outperforms the baseline methods based on various evaluation metrics. Additionally, the model efficiency analysis highlights that route predictions can be generated in less than 500ms per 10k requests, largely thanks to the Spanning Route algorithm, which validates the suitability of RouteKG for real-time traffic applications. To demonstrate the applicability of RouteKG beyond routing tasks, we utilize it to estimate link-level traffic flows, achieving an $\mathrm{R}^{2}$ value of 0.997 under the Goal scenario. This could provide valuable insights for future designs of intelligent transportation systems.
@@ VI-A Minibatch version of the Spanning Route algorithm
road network $\mathbf{G}=(\mathbf{V},\mathbf{E})$; NAE matrix $\mathbf{A}\in\mathbb{R}^{|\mathbf{V}|\times N_{A}}$; tree’s degree $n$.
@@
3// Initialize the root node.
@@
13             // Get the top-$n$ adjacent edges with highest probabilities based on $\mathrm{Pr}(\widetilde{x^{f,\gamma}_{i}})$.
14             $\left\{e_{k}^{\Gamma+\gamma}\in\mathbb{R}^{\mathcal{B}}\right\}_{k=1}^{n}$ = GetTopK($\mathrm{Pr}(\widetilde{x^{f,\gamma}})[:,\mathcal{N}_{end\_node}^{e}]$, $K=n$)
@@
18                   node = CreateNewNode(name=“$k$”, parent=leaf, end_node=$e_{k}^{\Gamma+\gamma}[1]\in\mathbb{R}^{\mathcal{B}}$, pred=$e_{k}^{\Gamma+\gamma}\in\mathbb{R}^{\mathcal{B}}$)
19
@@
21leaves $\leftarrow$ GetLeaves(root)
@@ VI-B Hyperparameters
This study applied consistent experimental configurations to both datasets to ensure reliable and comparable results. A training batch size of 2048 was maintained, and the maximum number of epochs was set to 10,000. Early stopping was employed with a patience parameter of 100 epochs. The hidden dimensions, denoted by $\delta_{\mathcal{E}}$, $\delta_{\mathcal{R}^{c}}$, $\delta_{\mathcal{R}^{s}}$, $\delta_{\mathcal{R}^{a}}$, and $\delta_{\mathcal{R}^{d}}$, were uniformly set to 64.
The Adam optimizer was utilized to update the model’s parameters, with a learning rate 1e-3 and weight decay set at 1e-2. To scale the sampling probability distribution for top-$k$ predictions, a temperature parameter of 0.1 was employed for the Chengdu dataset and 0.13 for the Shanghai dataset.
Regarding the weights assigned to different loss terms, under the NoGoal scenario, the weights [$w_{rep}$, $w_{rank}$, $w_{pred}$, $w_{d}$] were set at $[1,1,1,2.4]$ for the Chengdu dataset and $[1.3,2.8,0.5,2.9]$ for the Shanghai dataset. Under the GoalD scenario, the weights $[w_{rep},w_{rank},w_{pred}]$ were established as $[1,1,1]$ for the Chengdu dataset and $[1.4,2.1,1.7]$ for the Shanghai dataset. Lastly, for the Goal scenario, the weights $[w_{rep},w_{rank},w_{pred}]$ were set as $[2.4,2.2,2.8]$ for the Chengdu dataset and $[1.9,1.3,2.4]$ for the Shanghai dataset. These weight configurations were chosen based on preliminary experiments and empirical observations.

### v4

@@ V-C2 Main Results Analysis
Table III shows a comparison of the accuracy of the different methods in predicting future routes on two real-world datasets under three scenarios. Overall, RouteKG consistently outperforms all baselines across all evaluation metrics. It is observed that for all models, the route-level prediction accuracy is lower than the link-level prediction accuracy. This underscores the importance of modeling the consistency between different road links. Comparing the different models, we find that deep learning methods achieve higher accuracy in general and can be further enhanced by integrating additional information. For instance, models like NetTraj and RouteKG, which incorporate spatial data, outperform simpler models like RNN and its variants. Remarkably, RouteKG outperforms the NetTraj model and other baselines, even without extra information, which highlights the effectiveness of our approach to integrate KG for future route prediction. Comparing under different experimental settings, intuitively, introducing more goal information progressively improves overall accuracy. In particular, RouteKG’s prediction accuracy is greatly improved after the gradual incorporation of goal information, and at the same time, it also has an accuracy improvement of about 5.41% in comparison with the optimal baseline under the NoGoal condition, which proves the effectiveness of RouteKG in utilizing goal information and its applicability under various conditions. Lastly, comparing different datasets, the prediction accuracy of the Chengdu dataset is better than that of the Shanghai dataset, likely because of the larger data size of the former.
@@ V-D Ablation Analysis
Figure 8 compares the performance of RouteKG with its two ablation variants, where L-Rec denotes link-level recall, R-Rec denotes route-level recall, and R-MRR stands for route-level mean reciprocal rank. Notably, RouteKG w/o rerank removes the Rank Refinement Module. Experimental results show that removing this module significantly reduces prediction performance. This suggests the interconnected nature of link choices, emphasizing the need for a module to model route consistency and choice correlation. This highlights the module’s indispensability. Remarkably, even without reranking, RouteKG still outperforms most benchmark methods, particularly in the top 5 and top 10 predictions. This demonstrates RouteKG’s efficacy in identifying potential future routes, reinforcing the importance of integrating the ranking refinement module for enhancing top-1 predictions. The detailed ablation analysis results are shown in Appendix VI-C.
@@
TABLE IV: Ablation on relation types (NoGoal, Chengdu, link-level R@1).
Setting R@1 $\Delta$
All relations 0.841
   w/o DirectionTo 0.820 $2.1\%$
   w/o ConsistentWith 0.824 $1.7\%$
   w/o ConnectBy 0.831 $1.0\%$
   w/o DistanceTo 0.832 $0.9\%$
We further conduct ablation experiments on individual relation types to understand their respective contributions. As shown in Table IV, removing any single relation leads to a modest drop in performance (within 2% in link-level R@1), with DirectionTo being the most influential. These results suggest that while the model can learn effectively through the KGC formulation, KG representation learning still provides complementary relational priors that enhance prediction, especially under sparse supervision scenarios.
@@ V-E Sensitivity Analysis
This section presents a sensitivity analysis to assess the robustness and reliability of our model under various parameter settings. We first investigate the model performance under varying lengths of future routes to be predicted, denoted as $\Gamma^{\prime}$. As depicted in Figure 9, by altering $\Gamma^{\prime}$ from 2 to 8, there is a noticeable trend of declining performance with increasing $\Gamma^{\prime}$ values. This indicates that predicting longer routes becomes progressively challenging due to an expanded candidate space and uncertainty, particularly when lacking goal information.
@@
TABLE V: Performance of $\tau$-Pruned and 1-Greedy decoding strategies in Chengdu. Both maintain high accuracy with reduced computational cost.
Link-level Route-level
R@1 R@5 R@10 R@1 R@5 R@10 M@1 M@5 M@10
NoGoal 0.841 0.940 0.968 0.696 0.885 0.931 0.696 0.762 0.768
$\tau$-Pruned 0.830 0.935 0.965 0.690 0.880 0.925 0.690 0.758 0.765
1-Greedy 0.821 - - 0.674 - - 0.674 - -
NetTraj 0.809 0.874 0.909 0.662 0.836 0.882 0.662 0.730 0.735
GoalD 0.916 0.978 0.988 0.815 0.953 0.974 0.815 0.866 0.869
$\tau$-Pruned 0.905 0.970 0.987 0.810 0.950 0.972 0.810 0.860 0.868
1-Greedy 0.898 - - 0.794 - - 0.794 - -
NetTraj 0.868 0.923 0.949 0.741 0.896 0.929 0.741 0.802 0.806
Goal 0.974 0.991 0.995 0.958 0.983 0.988 0.958 0.967 0.968
$\tau$-Pruned 0.965 0.990 0.994 0.955 0.980 0.986 0.955 0.965 0.966
1-Greedy 0.952 - - 0.948 - - 0.948 - -
NetTraj 0.876 0.918 0.941 0.782 0.894 0.923 0.782 0.825 0.829
We further evaluate two heuristic decoding strategies to understand their trade-offs in accuracy and computational efficiency: (1) $\tau$-Pruned retains only the next-step candidates whose transition probability exceeds a pruning threshold $\tau$ (set as 0.2). (2) 1-Greedy is a fully greedy decoding strategy that selects only the most probable next-step link at each prediction step, effectively reducing the candidate set size to one per step.
Table V summarizes the performance of these variants on the Chengdu dataset. While both heuristics lead to slight reductions in performance compared to the full search, they still maintain competitive accuracy. For example, $\tau$-Pruned achieves a route-level R@10 of 0.986 (vs. 0.988 in full search), and 1-Greedy reaches 0.948 with goal information. These results highlight that RouteKG retains strong predictive performance even with aggressive pruning, validating the model’s robustness and generalization under restricted search space.
@@ V-F Efficiency Analysis
Efficient route prediction in transportation systems is crucial for ensuring prompt responses for system operators and road users. To assess the model efficiency, we analyze the inference time of various models in two datasets. Figure 10 delineates the inference times across various models. Note that the results from Dijkstra and RCM-BC are omitted due to their overly long inference times. All baselines utilize the Spanning Route algorithm to generate future routes except the Markov model, which uses pre-computed transition probabilities to sample and generate the top-$k$ predictions through $k$ iterations.
Figure 10: Left two: Inference time per 10K requests for different methods in Chengdu and Shanghai (mean ± std). Right: Inference time vs. rollout horizons, iterative top-1 rollouts show linear runtime growth.
TABLE VI: Traffic flow estimation results on Chengdu and Shanghai datasets. (mean ± std)
@@
RouteKG demonstrates remarkable efficiency, achieving average inference times of 598.01ms and 244.47ms for every 10k requests on the Chengdu and Shanghai datasets, respectively, with standard deviations of 1.21ms and 19.35ms. In contrast, the Dijkstra model, based on dynamic programming, takes over 38s, and the RCM-BC model exceeds 1000s, rendering them impractical for real-time systems. Models utilizing the Spanning Route algorithm (e.g., RNN, GRU, LSTM, NetTraj, RouteKG) show superior inference times, with less than 400ms in Chengdu and 250ms in Shanghai per 10k requests. RouteKG exhibits a marginally higher inference time, likely attributable to the reranking process.
Note that our framework is designed for short-term route prediction, where local transition patterns remain stable. For longer prediction horizons, directly increasing the number of steps is impractical, as it leads to exponential growth in candidate routes under the Spanning Route framework. We discuss these limitations and possible extensions in the Discussion section. To support longer predictions efficiently, we adopt a recursive rollout strategy. In this setting, the model predicts a short future segment (e.g., 5 steps) and recursively appends the top-1 predicted links as input for the next segment, enabling longer predictions without modifying the base architecture. This strategy leverages RouteKG’s high top-1 accuracy, making it suitable for multi-step rollout.
As shown in the right panel of Figure 10, inference time increases approximately linearly with the number of rollout iterations, under the same computational resources. This confirms that our framework remains computationally efficient even as the prediction horizon extends. Unlike conventional methods, where prediction cost grows exponentially due to expanding search trees, our strategy avoids full enumeration by committing to a single most likely route per segment. This makes the model suitable for time-sensitive applications where latency constraints prohibit full route expansion.
@@ V-G Case Study: Traffic Flow Estimation
In this section, to demonstrate the practical use cases of RouteKG, we conduct a case study on traffic flow estimation utilizing the model’s capability to generate future routes, which can then support other traffic control or management strategies. Specifically, we adopt a sampling-based method for generating future routes to maximize the utility of the top-$k$ future route predictions. Initially, the top-$k$ predictions are converted to a probability distribution using temperature scaling [77]. Subsequently, we sample from the predicted top-$k$ future routes for each observed trajectory based on their probability distribution. The estimated link-level traffic flows are then obtained by aggregating the number of predicted future routes. To ensure robustness, we iterate the experiments ten times and report traffic flow estimation results in a mean±std format, focusing solely on the top-10 predictions for simplicity.
The effectiveness of traffic flow estimation with RouteKG is demonstrated using three standard regression metrics: Mean Absolute Error (MAE), Root Mean Squared Error (RMSE), and the coefficient of determination ($\mathrm{R}^{2}$), as detailed in Table VI. RouteKG consistently outperforms in all metrics for both datasets, aligning with our main experiment results in Section V-C. As expected, incorporating more goal information leads to improved accuracy in traffic flow predictions.
In particular, RouteKG’s performance in the NoGoal scenario significantly surpasses the baseline for both datasets, suggesting that our method of estimating moving directions and leveraging KGC is more effective than current state-of-the-art (SOTA) modeling methods. Quantitatively, it reduces MAE, RMSE, and $\mathrm{R}^{2}$ by 34.7%, 43.5%, and 2.6%, respectively, compared to the best baseline. Under the GoalD scenario, performance increases notably, indicating potential for future refinement in modeling future directions. Importantly, RouteKG’s enhancements in traffic flow estimation, especially when including the actual future direction, are more significant than those of the baselines. This reaffirms RouteKG’s integration of direction information in the KGC problem. With actual goal information incorporated, RouteKG achieves an MAE of 1, RMSE of 3, and 99.7% in $\mathrm{R}^{2}$, underlining its efficacy and promise for practical applications.
@@
V-H Discussion
While the current framework builds on a static KG to model road network relations, it is designed with a focus on building a general and flexible solution based on the fundamental elements of road networks and could flexibly incorporate both static and dynamic contextual features. Real-world factors such as rush hours, weather, road closures, and special events can significantly affect route choices. Although these factors are not explicitly modeled in the current version, the framework naturally supports their integration. Contextual features can be embedded into the entity representations to reflect time-dependent behaviors or region-specific characteristics. For example, embeddings related to rush hours or nearby POIs can adjust the representation of road links to capture patterns such as congestion avoidance or attraction to popular destinations. These context-aware embeddings can be learned during KG representation learning and updated continuously as new data arrives, allowing the KG to evolve and better reflect realistic routing behavior.
Additionally, disruptions such as accidents or road closures can be handled by updating the NAE matrix, which dynamically masks affected links from the candidate space during prediction. This simple mechanism enables the model to adapt to real-time changes without retraining. RouteKG is also promising for sparse road networks, where simpler topology reduces prediction complexity. Even with fewer links, spatial relations in the KG can provide useful priors. While these extensions are not included in the current implementation, they can be incorporated into the existing framework with minimal modifications and are left for future work.
Finally, although our framework is optimized for short-term predictions, directly increasing the prediction length may cause rapid growth in the search space and inference time under the Spanning Route paradigm. We explore two heuristic strategies to mitigate this issue: pruning low-probability candidates and recursively rolling out top-1 predictions in short segments. These strategies reduce computational overhead and demonstrate promising performance in our sensitivity and efficiency studies. Nonetheless, efficiently scaling to longer prediction horizons remains an open challenge and a key direction for future improvement. In particular, enabling variable-length prediction, where the model learns to determine when to terminate based on the observed partial route (e.g., via a special “[End]” token), represents a promising avenue for future research.
@@ VI Conclusion
RouteKG is evaluated using taxi trajectory data from Chengdu and Shanghai. The evaluation considers three practical scenarios with different levels of goal information availability: NoGoal, GoalD, and Goal. The experiment results show that the proposed RouteKG consistently outperforms the baseline methods based on various evaluation metrics. Additionally, the model efficiency analysis highlights that route predictions can be generated in less than 500ms per 10k requests, largely thanks to the Spanning Route algorithm, which validates the suitability of RouteKG for real-time traffic applications. To demonstrate the applicability of RouteKG beyond routing tasks, we utilize it to estimate link-level traffic flows, achieving an $\mathrm{R}^{2}$ value of 0.997 under the Goal scenario.
@@
Acknowledgment
This research is supported by National Natural Science Foundation of China (42201502), and Seed Funding for Strategic Interdisciplinary Research Scheme at The University of Hong Kong (102010057).
@@ VI-A Minibatch version of the Spanning Route algorithm
road network $\mathbf{G}=(\mathbf{V},\mathbf{E})$;
NAE matrix $\mathbf{A}\in\mathbb{R}^{|\mathbf{V}|\times N_{A}}$; tree’s degree $n$.
@@
31ex
4// Initialize the root node.
@@
[…]
@@
[…]
@@
21
22leaves $\leftarrow$ GetLeaves(root)
@@ VI-B Hyperparameters
This study applied consistent experimental configurations to both datasets to ensure reliable and comparable results. All hyperparameters were selected by grid search on a held‐out validation set, ensuring that our final choices lie within the searched ranges. Below, we summarize both the search grids and the selected values.
We used a batch size of 2048 and trained for up to 10,000 epochs, with early stopping after 100 epochs without validation improvement. All embedding dimensions $\delta_{\mathcal{E}},\delta_{\mathcal{R}^{c}},\delta_{\mathcal{R}^{s}},\delta_{\mathcal{R}^{a}},\delta_{\mathcal{R}^{d}}$ were searched over $\{32,64,128\}$ and set to 64 in the final model. The Adam optimizer was used with weight decay chosen from $\{10^{-4},10^{-3},10^{-2}\}$; we selected $10^{-2}$. Learning rates were searched over $\{10^{-4},10^{-3},10^{-2}\}$ and fixed to $10^{-3}$. We tuned the sampling temperature from $\{0.05,0.1,0.15,0.2\}$. The best values were 0.10 for Chengdu and 0.13 for Shanghai. We balanced the four objectives in Eq. (16) by performing a greedy search over a predefined grid for each weight $w$ from 0.5 to 3. Starting from equal weights, we iteratively adjusted one component at a time based on validation performance (R@1), holding the others fixed, until no further improvement was observed. The final configurations selected from this grid are as follows. For the NoGoal setting, we used $w_{rep},w_{rank},w_{pred},w_{d}]=[1.0,1.0,1.0,2.4]$ on the Chengdu dataset, and $[1.3,2.8,0.5,2.9]$ on the Shanghai dataset. In the GoalD setting, the selected weights were $[1.0,1.0,1.0]$ for Chengdu and $[1.4,2.1,1.7]$ for Shanghai. For the Goal setting, we used $[2.4,2.2,2.8]$ and $[1.9,1.3,2.4]$ for Chengdu and Shanghai, respectively. All selected values lie within the predefined grid.
VI-C Numerical Results of Ablation Analysis
Table VII shows the detailed ablation analysis results under the GoalD setting.
TABLE VII: Full ablation analysis results under GoalD.
Chengdu Link-level Route-level
R@1 R@5 R@10 R@1 R@5 R@10 M@1 M@5 M@10
RouteKG 0.916 0.979 0.988 0.815 0.953 0.974 0.815 0.866 0.869
w/o rerank 0.874 0.934 0.960 0.732 0.910 0.943 0.732 0.804 0.808
w/o relation 0.910 0.979 0.989 0.805 0.953 0.973 0.805 0.861 0.864
Shanghai Link-level Route-level
R@1 R@5 R@10 R@1 R@5 R@10 M@1 M@5 M@10
RouteKG 0.843 0.946 0.963 0.723 0.894 0.918 0.723 0.780 0.784
w/o rerank 0.793 0.855 0.895 0.658 0.804 0.856 0.658 0.713 0.719
w/o relation 0.837 0.942 0.962 0.717 0.891 0.915 0.717 0.776 0.779